# RadGPT: Constructing 3D Image-Text Tumor Datasets

### v1

@@ 4.3 RadGPT Locates Tumors in Organs
RadGPT uses use organ sub-segments to locate tumors. It achieved a Dice similarity coefficient (DSC) of 0.85 in segmentingeightliver sub-segments, according to the test set from Zhang et al. [69]. For pancreas sub-segmentation, we do not have a ground-truth or dataset for testing, because AbdomenAtlas 3.0 is the first public dataset to present pancreas sub-segments (head, body, and tail). However, our algorithm to sub-segment the pancreas closely follows radiologist-accepted standards (Figure 4), and we asked radiologists to qualitatively evaluate our annotations.
@@
Figure 4: Pancreas and liver sub-segments. (a) RadGPT segments the pancreas based on radiology standards [58]. (I-II) the SMA separates the pancreas head (H) from the body (B), and (III) the remaining pancreas is divided at its midpoint into the body and tail (T). (b) Our liver sub-segmentation model achieved a DSC of 0.85 in segmentingeightliver sub-segments on a public test set [69]. Sub-segments are in different colors and tumors in brown. Sub-segments are essential for RadGPT to localize tumors.
4.4 RadGPT Enables Diagnostic Evaluation
Model BLEU METEOR ROUGE-1 R.-2 R.-L BERT
M3D 0.0 2.5 6.9 1.0 4.5 44.2
CT2Rep 0.0 4.4 7.7 0.7 5.3 51.3
RadGPT-S 1.0 11.0 14.8 1.8 8.9 54.9
RadGPT-N 0.7 11.2 24.7 2.8 12.4 53.4
Table 4: Report style impacts standard evaluation metrics. With the same dataset as Table 2, we evaluate the fully-automated reports from RadGPT with text similarity metrics. RadGPT narrative (N) and structured (S) reports differ in style only, having the same sensitivity and specificity for tumor detection. However, by mimicking the style of ground-truth reports (§3.2), narrative reports have 67.1% higher ROUGE-1.
In Table 4, we use standard text similarity metrics (common in LLM evaluation) to evaluate the reports generated by RadGPT, CT2Rep, and M3D. RadGPT achieves the highest scores in all metrics. These results align with the superiority of RadGPT in our diagnostic evaluation (Table 2). Thus, diagnostic accuracy may improve LLM metrics.
However, these metrics are also affected by the style of the report. Table 4 includes structured and narrative reports by RadGPT. They contain the same diagnoses and details, but different style (§3.2). In Table 4, ROUGE [39] was the metric most affected by the style variation and METEOR the least. The results show that standard LLM metrics are influenced by style, but the extent of this influence varies between metrics. Conversely, our proposed LLM-based sensitivity and specificity metrics are only influenced by the diagnostic accuracy of the reports. In addition, sensitivity and specificity provide clinicians with an objective and easily interpretable evaluation of AI-made reports, directly evaluating the clinical usefulness of a vision-language model.
4.5 RadGPT Can Enhance Human-made Reports
Human-made reports often have inconsistent documentation of critical findings, complicating clinical decision-making by referring physicians [1]. To improve consistency and interpretability, experts advocate for structured reports. AbdomenAtlas 3.0 and RadGPT address this human report problem providing structured and narrative reports.
Both types of reports provide measurements of organ and tumor volume and attenuation, rarely found in human-made reports. Of 90 randomly selected UCSF human-made reports, none had volumes or HU values. Volumetric assessments better capture tumor mass, detect size changes earlier, and improve prognostic accuracy [63, 46, 11]. Organ volume and attenuation for kidney, spleen, pancreas, and liver are valuable for monitoring and diagnosing diseases. E.g., reduced pancreatic volume can indicate diabetes [45] and low liver attenuation can indicate fat infiltration [49]. RadGPT and AbdomenAtlas 3.0 include volumes and HU values for organs and tumors. With time constraints, radiologists typically measure only the largest tumors in a CT. Of 90 human-made tumor reports from UCSF, 33 (36.7%) missed size or location of one or more tumor. Conversely, AbdomenAtlas 3.0 and RadGPT provide precise measurements and locations for all detected tumors. In summary, our structured reports have details usually absent in human-made reports: locations and measurements for all detected tumors, and tumor/organ volumes and HU values.
Human-AI fusion reports enhance human-made reports or clinical notes with the precision and detail of RadGPT structured reports. AbdomenAtlas 3.0 has clinical notes for 240 CT scans (209 are public, for TotalSegmentator CT scans [4]). However, these notes lack quantitative details in reports from RadGPT: of the 63 TotalSegmentator notes that mention tumors, none includes measurements, and many mention diverse tumors but do not describe them individually. However, these clinical notes describe diagnoses beyond the scope of RadGPT (which focuses on tumor-related diagnoses), such as calcified arterial plaques. We used an LLM to merge clinical notes with detailed structured and narrative reports in AbdomenAtlas 3.0, resulting in 240 fusion reports that integrate general clinical findings with precise tumor data (see Figure 11).
@@ 5 Conclusion & Future Work
AbdomenAtlas 3.0 is the first public dataset providing high-quality abdominal CT scans with reports and per-voxel tumor annotations, encompassing 9,262 CT scans from 88 institutions. It uniquely includes pancreas sub-segments, peripancreatic blood vessels, and pancreatic cancer stages, features absent in existing public datasets. RadGPT is an Anatomy-Aware Vision-Language AI Agent that transforms per-voxel annotations into structured reports using deterministic algorithms. These reports align with the accuracy of pathology findings annotated by a team of radiologists in AbdomenAtlas 3.0. Additionally, RadGPT enables fully automated report generation, surpassing existing approaches in detecting tumors from abdominal CT scans. Together, AbdomenAtlas 3.0 and RadGPT bridge the gap between tumor segmentation and report generation, offering valuable resources and tools to advance AI in automated abdominal CT interpretation.
We are committed to expanding AbdomenAtlas 3.0 to include reports for more types of tumors. Additionally, we plan to host benchmarks using AbdomenAtlas 3.0 with two train/test splits. IID Split: Randomly sets aside 10% of the dataset for testing, where training and testing data come from the same institutions, following standard AI evaluation practices. OOD Split: Uses data from 23 unseen institutions (4,500 CT scans) for testing, providing a large test set to evaluate AI generalization to new environments. This benchmark will assess report generation models using standard text similarity metrics but will prioritize tumor detection sensitivity and specificity, enhanced by our proposed LLM-based diagnostic evaluation.
Acknowledgments. This work was supported by the Lustgarten Foundation for Pancreatic Cancer Research and the Patrick J. McGovern Foundation Award. We thank the funding of Italian Institute of Technology and the HPC infrastructure at Italian Institute of Technology.
@@ A.1.2 Pancreas Sub-Segments
Figure 7: CT scan with 2 pancreatic tumors (yellow), and illustration of pancreas sub-segmentation into head (white, left), body (yellow, middle) and tail (white, right). RadGPT used the sub-segments to localize both PDAC tumors in the pancreas head. AbdomenAtlas 3.0 is the first to present pancreas sub-segments annotated per voxel. This information is crucial for writing radiology reports, as localizing pancreatic tumors in the pancreas head, body or tail is key for determining if the tumor can be surgically removed, and for tracking tumors in time.
Figure 8: CT scan with a pancreatic cyst (yellow), and illustration of pancreas sub-segmentation into head (white, left), body (yellow, middle) and tail (white, right). RadGPT used the sub-segments to localize the cyst in the pancreas head. AbdomenAtlas 3.0 is the first to present pancreas sub-segments annotated per voxel. This information is crucial for writing radiology reports, as localizing pancreatic tumors in the pancreas head, body or tail is key for determining if the tumor can be surgically removed, and for tracking tumors in time.
@@ A.1.4 Kidney Tumor Report
A.1.5 Human-AI Fusion Reports
Figure 11: In our human-AI fusion reports, LLMs combine detailed quantitative data from RadGPT’s reports with the generality of human-made reports or clinical notes. In the image, the LLM extracted kidney tumor sizes and volumes from the RadGPT report, while incorporating tumor type and non-cancer-related findings from the radiologist’s notes. Reports created by RadGPT include detailed information often absent in human-made reports, such as tumor volumes and individual measurements for a large number of tumors when present. In contrast, human-made reports provide general findings that can expand RadGPT reports to many insights beyond cancer.
@@ A.1.6 Narrative Reports
Figure 12: Example of Narrative Report: we use LLM to convert structured reports into narrative reports that follow the writing style of a target hospital’s. The LLM receives 10 example reports from the hospitals as examples of style, and is instructed not to change the medical content of the structured report during style adaptation. Since reports targeting diverse abnormalities vary strongly in style, working ans structure, we use another LLM to pre-classify the hospital’s human-made reports into diagnostic categories (e.g., liver tumor). Thus, during style adaptation, we use as examples only reports that focus on the same diagnosis as the structured report. E.g., if the structured report mentions liver tumors, the examples also will concentrate on liver tumors.
@@ Appendix B Technical Details of RadGPT
2:  for each CT slice $s$ containing tumor $A$ do
@@
11:  Report the tumor size as $D_{max}\times d$, converting from mm to cm.
@@
30:     Determine stage by tumor size: T1a $\leq$ 0.5cm, T1b 0.5–1cm, T1c 1–2cm, T2 2–4cm, T3 $>$ 4cm
@@ B.2 Segmentation Post-processing
Figure 14: Tumor detection sensitivity and specificity for RadGPT with diverse thresholds. Evaluation performed on a private dataset from a hospital never observed during training, detailed in Table 2.
B.3 LLM Prompts
B.3.1 Style Adaptation
@@
[…]
B.3.2 AI-Human Report Fusion
@@
[…]
B.3.3 Labeling/Report Evaluation
@@
B.4 Organ size standards
@@
Appendix C Detailed Tumor Statistics

### v2

@@ 4.2 RadGPT Locates Tumors in Organs
RadGPT uses use organ sub-segments to locate tumors. It achieved a Dice similarity coefficient (DSC) of 0.85 in segmenting eight liver sub-segments, according to the test set from Zhang et al[58]. For pancreas sub-segmentation, we do not have a ground-truth or dataset for testing, because AbdomenAtlas 3.0 is the first public dataset to present pancreas sub-segments (head, body, and tail). However, our algorithm to sub-segment the pancreas closely follows radiologist-accepted standards (see Figure 5), and we asked radiologists to qualitatively evaluate our annotations.
@@
Figure 5: Pancreas and liver sub-segments. (a) RadGPT segments the pancreas based on radiology standards [48]. (I-II) the SMA separates the pancreas head (H) from the body (B), and (III) the remaining pancreas is divided at its midpoint into the body and tail (T). (b) Our liver sub-segmentation model achieved a DSC of 0.85 in segmenting eight liver sub-segments on a public test set [58]. Sub-segments are in different colors and tumors in brown. Sub-segments are essential for RadGPT to locate tumors.
4.3 RadGPT Enhances Human-made Reports
Human-made radiology reports often omit critical quantitative details—such as tumor volumes and attenuation (HU) values—compromising clinical decision-making. In our evaluation of 90 human reports from a UCSF, none reported organ or tumor volumes and only 63% measured all detected tumors. In contrast, our structured reports (RadGPT) consistently provide full, quantitative data. As shown in Table 4, while human-made reports measure volume and HU in 0% of cases and capture all tumors in only 63% of cases, RadGPT achieves 100% for all metrics. This level of consistency streamlines clinical assessments and improves prognostic accuracy by ensuring every tumor is precisely measured [53, 39, 10].
In AbdomenAtlas 3.0, 240 CT scans include clinical notes, which lack quantitative tumor measurements; for instance, among 63 TotalSegmentator notes mentioning tumors, none provide such data—even though they report other findings like calcified arterial plaques. By merging these notes with our structured and narrative reports using an LLM, RadGPT generates 240 enhanced reports that integrate the notes’ comprehensive clinical findings (covering 66 diagnoses) with precise tumor sizes (see Figure 12).
4.4 Discussion: End-to-End, Segmentation-based?
Table 2 shows RadGPT outperformed end-to-end VLMs. The unique design of RadGPT offers several advantages over end-to-end training. (i) Interpretable: RadGPT generates reports from tumor segmentation, allowing clinicians and developers to visualize and verify tumor locations and sizes in the CT. In contrast, errors in end-to-end methods are harder to diagnose and debug. (ii) Interactive: Our goal is creating a high-quality dataset of CTs and reports to drive innovation in report generation. As all algorithms make mistakes, a human-in-the-loop approach is key. Segmentation is a safeguard, allowing radiologists to ensure reports are correct, and easily transforming radiologist-revised masks into reports. (iii) Strong supervision: tumor segmentation AI has been a long-term focus of the research community, achieving high accuracy by leveraging precise per-voxel masks. Our benchmark shows that segmentation-assisted models can transfer this high accuracy to reports.
Volume HU Diameters
Human-made 0% 0% 63%
RadGPT (ours) 100% 100% 100%
Table 4: Comparison of human-made reports vs. RadGPT reports for 90 UCSF CTs. Values indicate the percentage of reports containing tumor volume, HU and diameter measurements for all detected tumors. RadGPT reports provide more clinically relevant [53, 39, 10] quantitative details about tumors.
@@ 5 Conclusion & Future Work
Dataset curation and report‑generation are inter‑dependent. Developing image-report-mask methods requires image-report-mask datasets, but creating these datasets requires reliable methods and human-in-the-loop involvement. Our focus is creating high-quality image-report-mask datasets to support further methodological advancements.
AbdomenAtlas 3.0 is the first public dataset providing high-quality abdominal CT scans with reports and per-voxel tumor annotations, encompassing 9,262 CT scans from 138 institutions. It uniquely includes pancreas sub-segments, peripancreatic blood vessels, and pancreatic cancer stages—absent in existing public datasets. RadGPT transforms per-voxel annotations into structured reports using deterministic algorithms. These reports align with the accuracy of segmentations revised by radiologists in AbdomenAtlas 3.0. Additionally, RadGPT enables fully-automated report generation, surpassing existing approaches in detecting tumors. Together, AbdomenAtlas 3.0 and RadGPT bridge the gap between tumor segmentation and report generation, offering valuable resources and tools to advance AI in abdominal CT interpretation.
We are committed to expanding AbdomenAtlas 3.0 to include reports for more types of tumors. Additionally, we plan to host benchmarks using AbdomenAtlas 3.0 with two train/test splits. IID Split: Randomly sets aside 10% of the dataset for testing, where training and testing data come from the same institutions, following standard AI evaluation practices. Used in Table 2. OOD Split: Uses data from 23 unseen institutions (4,500 CT scans) for testing, providing a large test set to evaluate AI generalization to new environments. This benchmark will assess report generation models using standard text similarity metrics but will prioritize tumor detection sensitivity and specificity, enabled by our proposed LLM-based diagnostic evaluation.
Although 66 out of 240 fusion reports present diverse diagnoses, AbdomenAtlas 3.0 is cancer-centric. Cancer is a leading cause of death, and over 40% of medical imaging reports focus on cancer detection. Thus, AI-assisted report generation has the potential for significant impact. We hope our release of the cancer-centric AbdomenAtlas 3.0 will stimulate further AI advancements in the field.
Acknowledgments. This work was supported by the Lustgarten Foundation for Pancreatic Cancer Research and the National Institutes of Health (NIH) under Award Number R01EB037669. We would like to thank the Johns Hopkins Research IT team in IT@JH for their support and infrastructure resources where some of these analyses were conducted; especially DISCOVERY HPC. We thank the funding of Italian Institute of Technology and the HPC infrastructure at Italian Institute of Technology. P.R.A.S.B. thanks the funding from the Center for Biomolecular Nanotechnologies, Istituto Italiano di Tecnologia (73010, Arnesano, LE, Italy).
@@ A.1.2 Pancreas Sub-Segments
Figure 8: CT scan with 2 pancreatic tumors (yellow), and illustration of pancreas sub-segmentation into head (white, left), body (yellow, middle) and tail (white, right). RadGPT used the sub-segments to locate both PDAC tumors in the pancreas head. AbdomenAtlas 3.0 is the first to present pancreas sub-segments annotated per voxel. This information is crucial for writing radiology reports, as localizing pancreatic tumors in the pancreas head, body or tail is key for determining if the tumor can be surgically removed, and for tracking tumors in time.
Figure 9: CT scan with a pancreatic cyst (yellow), and illustration of pancreas sub-segmentation into head (white, left), body (yellow, middle) and tail (white, right). RadGPT used the sub-segments to locate the cyst in the pancreas head. AbdomenAtlas 3.0 is the first to present pancreas sub-segments annotated per voxel. This information is crucial for writing radiology reports, as localizing pancreatic tumors in the pancreas head, body or tail is key for determining if the tumor can be surgically removed, and for tracking tumors in time.
@@ A.1.4 Kidney Tumor Report
A.1.5 Enhanced Human Reports
Figure 12: In our enhanced human reports, LLMs combine detailed quantitative data from RadGPT’s reports with the generality of human-made reports or clinical notes. In the image, the LLM extracted kidney tumor sizes and volumes from the RadGPT report, while incorporating tumor type and non-cancer-related findings from the radiologist’s notes. Reports created by RadGPT include detailed information often absent in human-made reports, such as tumor volumes and individual measurements for a large number of tumors when present. In contrast, human-made reports provide general findings that can expand RadGPT reports to many insights beyond cancer.
@@ A.1.6 Narrative Reports
Figure 13: Example of Narrative Report: we use LLM to convert structured reports into narrative reports that follow the writing style of a target hospital’s. The LLM receives 10 example reports from the hospitals as examples of style, and is instructed not to change the medical content of the structured report during style adaptation. Since reports targeting diverse abnormalities vary strongly in style, working and structure, we use another LLM to pre-classify the hospital’s human-made reports into diagnostic categories (e.g., liver tumor). Thus, during style adaptation, we use as examples only reports that focus on the same diagnosis as the structured report. E.g., if the structured report mentions liver tumors, the examples also will concentrate on liver tumors.
@@ Appendix B Technical Details of RadGPT
2:for each CT slice $s$ containing tumor AA do
@@
11: Report the tumor size as $D_{max}\times d$, converting from $m$ to cm.
@@
30:  Determine stage by tumor size: T1a $\leq$ 0.5cm, T1b 0.5–1cm, T1c 1–2cm, T2 2–4cm, T3 >> 4cm
@@
For CT-CHAT, we similarly trained the model using only CT scans and structured reports without using per-voxel annotations from AbdomenAtlas 3.0. While our general training approach again mirrored the original authors’ published code and hyper-parameters [20], specific adaptations included standardizing the CT spacing to an isotropic 1.5 x 1.5 x 1.5 mm resolution, selecting a patch size of 20 in each dimension (x, y, and z), and ensuring consistency by center-cropping or padding scans to uniform dimensions of 300 × 300 × 600 mm. Training was performed using four A100 GPUs for 20,000 iterations with a batch size of 16. Furthermore, we employed visual instruction fine-tuning identical to the CT-CHAT authors, using an attention pooling mechanism that reduced tokens generated by CT-CLIP to 256 via learned queries, which were then linearly transformed to match the hidden dimension of the Llama 3.1 8B model. Visual instruction fine-tuning proceeded for 100 epochs.
The training of Merlin also leveraged only CT scans and structured reports, again without incorporating per-voxel annotations from AbdomenAtlas 3.0. While closely aligning with the authors’ original published code and hyper-parameters [8], Merlin required a distinct approach since only the pretrained volume encoder optimized for the abdominal region was available, without the report-generation weights. Consequently, we conducted visual instruction fine-tuning by applying a linear transformation to Merlin’s encoded embeddings, mapping them to the hidden dimension of the Rad LLama2 7B model. This fine-tuning stage continued for 100 epochs, effectively adapting Merlin for abdominal CT report generation.
@@ B.2 Segmentation Post-processing
Figure 15: Tumor detection sensitivity and specificity for RadGPT with diverse thresholds. Evaluation performed on a private dataset from a hospital never observed during training (UCSF), detailed in Table 2.
B.3 RadGPT Enables Diagnostic Evaluation
model BLEU METEOR ROUGE-1 R.-2 R.-L BERT RadGraph-F1
CT2Rep 0.26 8.27 12.73 0.95 7.46 45.19 3.01
CT-CHAT 0.15 7.19 10.90 0.54 6.00 45.63 5.69
M3D 0.02 5.05 12.51 1.80 7.32 44.05 1.69
Merlin 0.17 7.30 11.63 0.72 6.38 45.44 6.20
RadFM 0.10 6.90 16.14 2.41 9.26 46.40 2.36
RadGPT-S 0.72 8.82 13.56 1.22 7.64 45.62 4.15
RadGPT-N 0.74 10.47 22.77 2.64 11.89 45.41 2.92
Table 5: Report style impacts standard evaluation metrics. Testing on the same unseen hospital as in Table 2 (UCSF), we evaluate the fully-automated reports from RadGPT with text similarity metrics and RadGraph-F1[56] (using RadGraph-XL [16], which covers the abdomen region). RadGPT narrative (N) and structured (S) reports differ in style only, having the same sensitivity and specificity for tumor detection. However, by mimicking just the style of the test hospital (UCSF, §3.2), our narrative reports (RadGPT-N) achieved considerably higher METEOR and ROUGE scores. In contrast, our proposed evaluation metric (Table 2) only evaluates the diagnosis in the reports and it is not sensible to style variations.
In Table 5, we use standard text similarity metrics (common in LLM evaluation) and RadGraph-F1 to evaluate the reports generated by the AI. RadGPT achieves the highest scores in BLEU, METEOR, and ROUGE. These results align with the superiority of RadGPT in our diagnostic evaluation (Table 2). Thus, diagnostic accuracy may improve LLM metrics. However, BERT and RadGraph-F1 are not aligned with the diagnostic accuracy in Table 2—RadGPT has considerably superior diagnostic accuracy for cancer (Table 2), but it does not have the highest BERT and RadGraph-F1 scores. Moreover, NLP metrics are affected by the style of the report. Table 5 includes structured and narrative reports by RadGPT. They contain the same diagnoses and details, but different style (§3.2), with the narrative reports mimicking the style of the test hospital. In Table 5, ROUGE [35] was the metric most affected by the style variation. The results show that standard LLM metrics are influenced by style, but the extent of this influence varies between metrics. Conversely, our proposed LLM-based sensitivity and specificity metrics are only influenced by the diagnostic accuracy of the reports. In addition, sensitivity and specificity provide clinicians with an objective and easily interpretable evaluation of AI-made reports, objectively measuring the clinical usefulness of a vision-language model.
B.4 LLM Prompts
B.4.1 Style Adaptation
@@
Example Reports (Target Style): $\{\text{examples}\}$
Structured Report to Paraphrase: $\{\text{structured\_report}\}$
B.4.2 Enhancing Human Reports
@@
$\{\text{structured\_report}\}$
B.4.3 Labeling/Report Evaluation
@@
B.5 Organ size standards
@@
Appendix C Revisions by Radiologists
In AbdomenAtlas 3.0, organ segmentation masks were created and verified by radiologists through an efficient human-in-the-loop approach [31]. Conversely, the tumor segmentation masks were suggested by AI and radiologists individually verified and corrected them. To verify reports efficiently, radiologists first ensured that structured reports correctly described the already revised per-voxel tumor annotations. This confirmation was key to ensure our deterministic algorithms worked correctly. Then, to verify our narrative and enhanced human reports, we first used our double-check procedure: an LLM (Llama 3.1 70B) extracted tumor information from the narrative / human enhanced reports and checked if it matched the information in the corresponding structured reports. For cases of mismatch, we prompted the LLM to correct the narrative / human enhanced reports and repeated the double-check. Any remaining mismatch was sent to radiologists. Mismatches were also analyzed to improve our prompts. E.g., radiologists identified that a few narrative or enhanced reports introduced findings absent from the source structured report or human report. To correct this, we started prompting the LLM to justify each finding by quoting sentences in the source structured report or human report. After the double check, we organized structured, narrative and enhanced human reports in a table and radiologists could quickly compare them, confirming they had consistent medical findings.
Appendix D Detailed Tumor Statistics